OpenStudy (anonymous):

How do I solve for the angle when cos2x+cosx=0? havent done in a while need refreshing

OpenStudy (anonymous):

No one? really need it

OpenStudy (anonymous):

I think it may be no solution?

OpenStudy (blacksteel):

There is one; I'm working on it

OpenStudy (anonymous):

okay thanks

OpenStudy (blacksteel):

If cos(x) + cos(2x) = 0, cos(x) = -cos(2x) We see 3 possible cases here. Since cos(k*pi) = 1 for even k and -1 for odd k, one solution is: x = pi + k*2*pi. Since 2x = 2*pi + 4*k*pi, one x will have an odd coefficient and 2x will have an even one, so cos(x) = -1 and cos(2x) = 1. The other two cases are when x = 1/3*pi + 2*pi*n (since cos(1/3*pi) = cos(2/3*pi) and adding 2*pi*n to 1/3*pi adds 4*pi*n to 2/3*pi, so the locations of the angles will be the same with). We see the exact same thing for -1/3*pi, so the other set of solutions is x = 2*pi*n - 1/3pi Thus: x = pi + 2*pi*n x = 1/3*pi + 2*pi*n x = -1/3*pi + 2*pi*n